OpenStudy (anonymous):

+1 medal +1 fan +1 fan testimonial Given the function f(x) = x2 and k = –3, which of the following represents a vertical shift... f(x) + k kf(x) f(x+k) f(kx)

OpenStudy (anonymous):

tititanc help me after

OpenStudy (anonymous):

Think about how the graph of y=x looks like compared to y=x+3. How about y=x^2 compared to y=x^2-7 In general, adding a constant to a function shifts it vertically. Choice b dilates the graph depending on the value of k. If k is a positive real number > 1 the graph will be stretched vertically (and thus compressed horizontally), for example. If k is a number < 1 but > 0, then the graph is compressed vertically and stretched horizontally (it becomes wider) Choice c moves the graph horizontally. Usually it is written as f(x-k), where the graph is moved k units. If k is positive, it's to the right, if it's negative (which will cause it to be f(x-(-k)) = f(x+k)) then it's to the left. Choice d dilates the graph too, but in the opposite way that b does. So if k is a positive real number > 1, then the graph will be dilated **horizontally** by k, but **vertically** by 1/k. In other words, it will be compressed/become wider like in choice b. If k is a number <1 but >0, then the graph is stretched vertically.